OpenStudy (linn99123):

A triangle has a base of 12 meters and a height of 9 meters. What is the area of the triangle? A. 21 square meters B. 42 square meters C. 54 square meters D. 108 square meters

OpenStudy (luigi0210):

\[\LARGE A=\frac{1}{2}bh\]

OpenStudy (linn99123):

i have this one aadn 3 more :D

OpenStudy (31356):

C

OpenStudy (31356):

12*9=108 divide by 1/2=54